# On point spectrum of Jacobi matrices generated by iterations of quadratic polynomials

### v1

@@ 2.3. Action of Ruelle operators on a class of rational functions
Let
$g_{z}(x)=\frac{A}{x-z},\quad z\in{\mathbb{C}}\setminus{\mathsf{E}}$
with a constant $2\times 2$ matrix $A$. Then
$({\mathcal{M}}_{\varkappa}g_{T^{\circ k}(0)})(x)=\frac{\Phi_{\varkappa}A\Phi_{{\varkappa}}^{*}}{-T^{\circ k}(0)(x-T(0))}+\frac{T^{\circ k}(0){\mathfrak{p}}_{\varkappa}(T^{\circ k}(0))A{\mathfrak{p}}_{\varkappa}(T^{\circ k}(0))^{*}}{x-T^{\circ(k+1)}(0)},$ (2.9)
@@
Let us apply ${\mathcal{M}}_{\varkappa}$ to $g_{z}(x)$. To this end we consider the function
$\frac{1}{2}{\mathfrak{p}}_{{\varkappa}}(z)\frac{AT^{\prime}(z)}{x-T(z)}{\mathfrak{p}}^{*}_{{\varkappa}}(z)=({\mathcal{M}}_{\varkappa}g_{z})(x)+\frac{B_{0}(x)}{z}$
@@
$B_{0}(x)=\frac{\Phi_{\varkappa}A\Phi^{*}_{\varkappa}}{x-T(0)}.$
Set $z=T^{\circ k}(0)$ and we obtain (2.9). ∎
@@
Define the Herglotz function $w^{n}_{x}(z)$ in $z$, $n\geq 0$, by
$w^{n}_{x}(z):=\frac{1}{2^{n}}\frac{(T^{\circ n})^{\prime}(z)}{x-T^{\circ n}(z)}=\frac{zT(z)\cdots T^{\circ(n-1)}(z)}{{x-T^{\circ n}(z)}},\quad x\in[-\xi,\xi],$
where by definition $T^{\circ 0}(z)=z$. We can rewrite (2.9) in terms of $w^{k}_{x}(T(0))$ as follows
@@
3. Main theorem
Note that
${\mathsf{E}}\subset[-\xi,\xi],\quad\xi=\frac{1+\sqrt{1+4\lambda}}{2}<\lambda,$
and these are exact upper and lower bounds for ${\mathsf{E}}$.
@@ 3. Main theorem
Let ${\varkappa}^{\prime}_{n}$ be a certain “extension” of ${\varkappa}$, that is, ${\varkappa}_{n}^{\prime}$ is such that
$({\mathfrak{M}}_{{\varkappa}_{n}^{\prime}}^{*}\Sigma_{{\varkappa}_{n}^{\prime}})(dx)=\Sigma_{{\varkappa}}(dx),\quad{\mathfrak{M}}_{{\varkappa}^{\prime}_{n}}={\mathcal{M}}_{\hat{\underline{s}}^{n-1}{\varkappa}_{n}^{\prime}}\cdots{\mathcal{M}}_{{\varkappa}_{n}^{\prime}}.$
@@
$h_{n}(x)=({\mathfrak{M}}_{{\varkappa}_{n}^{\prime}}h_{0})(x),\quad h_{0}(x)=w_{x}^{0}(T(0))I.$
Then the sequence $h_{n}(x)$ is uniformly bounded on $[-\xi,\xi]$.
@@
Acting one by one by the operators ${\mathcal{M}}_{\hat{\underline{s}}^{k}{\varkappa}_{n}^{\prime}}$ and applying (2.5) we obtain that $h_{n}(x)$ is of the form
$h_{n}(x)=\sum_{k=0}^{n}A_{k}^{n}w^{k}_{x}(T(0)),$
@@
Since
$w^{k}_{x}(T(0))=\frac{1}{2^{k}}\frac{(T^{\circ k})^{\prime}(T(0))}{x-T^{\circ k}(T(0))}=\frac{1}{2^{k}}\sum_{T^{\circ k}(y)=x}\frac{1}{y-T(0)}=\frac{1}{2^{k}}\sum_{T^{\circ k}(y)=x}\frac{1}{y+\lambda}$
and $\{y:T^{\circ k}(y)=x\}\subset[-\xi,\xi]$, we have
@@
$\int\text{\rm tr}(h_{n}(x)\Sigma_{{\varkappa}_{n}^{\prime}}(dx))=\int\text{\rm tr}(({\mathfrak{M}}_{{\varkappa}_{n}^{\prime}}h_{0})(x)\Sigma_{{\varkappa}_{n}^{\prime}}(dx))=\int\text{\rm tr}(h_{0}(x)\Sigma_{\varkappa}(dx))$
@@
$\int\text{\rm tr}(h_{0}(x)\Sigma_{\varkappa}(dx))=\int\text{\rm tr}(h_{n}(x)\Sigma_{{\varkappa}_{n}^{\prime}}(dx))\geq\frac{1}{\lambda+\xi}\sum_{k=0}^{n}\text{\rm tr}(A^{n}_{k}\int\Sigma_{{\varkappa}_{n}^{\prime}}(dx))$
@@
$\int\Sigma_{{\varkappa}_{n}^{\prime}}(dx)=I,$
@@
$\text{\rm tr}\,h_{n}(x)\leq\frac{\lambda+\xi}{\lambda-\xi}\int\text{\rm tr}(h_{0}(x)\Sigma_{\varkappa}(dx))$ (3.4)
Proof of the Main Theorem 1.1.
Assume that for some ${\varkappa}\in{\mathbb{Z}}_{2}$ and $x_{0}\in{\mathsf{E}}$
@@
$\int g(x)\text{\rm tr}\left((h_{n}(x)\Sigma_{{\varkappa}^{\prime}_{n}}(dx)\right)=$ $\int g(x)\text{\rm tr}\left(({\mathfrak{M}}_{{\varkappa}_{n}^{\prime}}h_{0})(x)\Sigma_{{\varkappa}^{\prime}_{n}}(dx)\right)$
$=$ $\int\text{\rm tr}\left(({\mathfrak{M}}_{{\varkappa}_{n}^{\prime}}h_{0}\cdot g\circ T^{\circ n})(x)\Sigma_{{\varkappa}_{n}^{\prime}}(dx)\right)$
$=$ $\int\text{\rm tr}\left(h_{0}(x)g(T^{\circ n}(x))({\mathfrak{M}}_{{\varkappa}_{n}^{\prime}}^{*}\Sigma_{{\varkappa}_{n}^{\prime}})(dx)\right)$
@@
We choose a subsequence $I_{x_{0}}$ such that
$\lim_{\begin{subarray}{c}n\to\infty\\ n\in I_{x_{0}}\end{subarray}}T^{\circ n}(x_{0})=y_{0}.$
Now, let us fix an arbitrary small ${\epsilon}$-vicinity of $y_{0}$, $V_{\epsilon}=({y_{0}-{\epsilon}},{y_{0}+{\epsilon}})$. For a sufficiently big $n\in I_{x_{0}}$ we have $T^{\circ n}(x_{0})\in V_{\epsilon}$ and threrefore
$\int_{V_{\epsilon}}\text{\rm tr}(h_{n}(x)\Sigma_{{\varkappa}_{n}^{\prime}}(dx))=\int_{(T^{\circ n})^{-1}V_{\epsilon}}\text{\rm tr}(h_{0}(x)\Sigma_{{\varkappa}}(dx))\geq\text{\rm tr}(h_{0}(x_{0})\Sigma_{{\varkappa}}(\{x_{0}\}))$
@@
$C\text{\rm tr}\Sigma_{{\varkappa}_{n}^{\prime}}(V_{\epsilon})\geq\text{\rm tr}(h_{0}(x_{0})\Sigma_{{\varkappa}}(\{x_{0}\}))$
with a uniform constant $C$.
Let ${\varkappa}_{n}^{\prime}=2^{n}{\varkappa}$. Passing to the limit in $n$ we get ${\varkappa}_{n}^{\prime}\to 0$ and
$C\text{\rm tr}\Sigma_{0}(V_{\epsilon})\geq\text{\rm tr}(h_{0}(x_{0})\Sigma_{{\varkappa}}(\{x_{0}\})).$
Since ${\epsilon}$ is arbitrary small, the limit measure $\Sigma_{0}$ associated to the initial $J_{0}$ has a non-trivial mass at $y_{0}$. But the spectrum of $J_{0}$ is singular continuous. We arrive to a contradiction, that is, $\Sigma_{{\varkappa}}(\{x_{0}\})=0$. ∎
Appendix A Some explicit formulas
A.1. Recursion for resolvent matrix functions
Lemma A.1.
Let
$W_{\varkappa}(z)=\int\frac{\Sigma_{\varkappa}(dx)}{x-z}$
be the resolvent matrix-function. The following recursion holds
$W_{\varkappa}(z)=-\frac{1}{z}\pi_{\varkappa}+z{\mathfrak{p}}_{\varkappa}(\bar{z})^{*}W_{\hat{}\underline{s}{\varkappa}}(T(z)){\mathfrak{p}}_{\varkappa}(z)$ (A.1)
where
$\pi_{\varkappa}=\begin{cases}\begin{pmatrix}1&0\\ 0&0\end{pmatrix},\quad{\varkappa}=2\underline{s}{\varkappa},\\ \\ \begin{pmatrix}0&0\\ 0&1\end{pmatrix},\quad{\varkappa}=1+2\underline{s}{\varkappa}.\end{cases}$
@@ A.1. Recursion for resolvent matrix functions
Recall that matrix orthogonal polynomials of the second kind are given by
$\int\frac{\Sigma(dx)}{x-z}\begin{pmatrix}p_{n}^{-}(x)-p_{n}^{-}(z)\\ -a_{0}(q_{n}^{-}(x)-q_{n}^{-}(z))\end{pmatrix}=\begin{pmatrix}q_{n}^{-}(z)\\ 0\end{pmatrix},$ (A.2)
$\int\frac{\Sigma(dx)}{x-z}\begin{pmatrix}-a_{0}(q_{n}^{+}(x)-q_{n}^{+}(z))\\ p_{n}^{+}(x)-p_{n}^{+}(z)\end{pmatrix}=\begin{pmatrix}0\\ q_{n}^{+}(z)\end{pmatrix}.$ (A.3)
According to this we define
${\mathfrak{q}}_{\varkappa}=\begin{cases}\begin{pmatrix}\frac{1}{a_{{\varkappa}-1}}&0\\ 0&0\end{pmatrix},\quad{\varkappa}=2\underline{s}{\varkappa},\\ \\ \begin{pmatrix}0&0\\ 0&\frac{1}{a_{{\varkappa}+1}}\end{pmatrix},\quad{\varkappa}=1+2\underline{s}{\varkappa}.\end{cases}$
Starting with (2.4), using orthogonality and (A.2)–(A.3), we have
$\frac{1}{2}\int\frac{T^{\prime}(z)}{x-T(z)}\Sigma_{\hat{}\underline{s}{\varkappa}}(dx)=\int\frac{{\mathfrak{p}}^{-1}_{\varkappa}(x)^{*}\Sigma_{{\varkappa}}(dx){\mathfrak{p}}_{\varkappa}(x)^{-1}}{x-z}$
$={\mathfrak{p}}^{-1}_{\varkappa}(\bar{z})^{*}\int\frac{\Sigma_{{\varkappa}}(dx){\mathfrak{p}}_{\varkappa}(x)^{-1}}{x-z}+\int\frac{({\mathfrak{p}}^{-1}_{\varkappa}(x)^{*}-{\mathfrak{p}}^{-1}_{\varkappa}(z)^{*})}{x-z}\Sigma_{{\varkappa}}(dx){\mathfrak{p}}_{\varkappa}(x)^{-1}$
$={\mathfrak{p}}^{-1}_{\varkappa}(\bar{z})^{*}\int\frac{\Sigma_{{\varkappa}}(dx)}{x-z}{\mathfrak{p}}^{-1}_{\varkappa}(z)+{\mathfrak{p}}^{-1}_{\varkappa}(\bar{z})^{*}\int\frac{\Sigma_{{\varkappa}}(dx)({\mathfrak{p}}_{\varkappa}(x)^{-1}-{\mathfrak{p}}_{\varkappa}(z)^{-1})}{x-z}$
$={\mathfrak{p}}^{-1}_{\varkappa}(\bar{z})^{*}\int\frac{\Sigma_{{\varkappa}}(dx)}{x-z}{\mathfrak{p}}^{-1}_{\varkappa}(z)+{\mathfrak{p}}^{-1}_{\varkappa}(\bar{z})^{*}{\mathfrak{q}}_{\varkappa}.$
We get (A.1). ∎
Lemma A.2.
Let ${\mathfrak{P}}^{n}_{\varkappa}(z)$ denote the product of matrix-functions
${\mathfrak{P}}^{n}_{\varkappa}(z)={\mathfrak{p}}_{\hat{}\underline{s}^{n}{\varkappa}}(T^{\circ n}(z))\cdots{\mathfrak{p}}_{\hat{}\underline{s}{\varkappa}}(T(z)){\mathfrak{p}}_{{\varkappa}}(z)$
Then
$1=w_{0}^{0}(T(0)){\langle\Psi_{\varkappa},\Upsilon_{\hat{\underline{s}}{\varkappa}}\rangle^{2}}+\sum_{k=2}^{\infty}w^{k-1}_{0}(T(0))\langle{\mathfrak{P}}^{k-2}_{\hat{\underline{s}}{\varkappa}}(T(0))\Psi_{\varkappa},\Upsilon_{\hat{}\underline{s}^{k}{\varkappa}}\rangle^{2}$ (A.4)
where
$\Psi_{\varkappa}=\begin{cases}\begin{pmatrix}a_{{\varkappa}-1}\\ a_{{\varkappa}}\end{pmatrix},\quad{\varkappa}=2\underline{s}{\varkappa},\\ \\ \begin{pmatrix}a_{{\varkappa}}\\ a_{{\varkappa}+1}\end{pmatrix},\quad{\varkappa}=1+2\underline{s}{\varkappa}.\end{cases}\quad\Upsilon_{\varkappa}^{*}=\begin{cases}\begin{pmatrix}1&0\end{pmatrix},\quad{\varkappa}=2\underline{s}{\varkappa},\\ \\ \begin{pmatrix}0&1\end{pmatrix},\quad{\varkappa}=1+2\underline{s}{\varkappa}.\end{cases}$

### v2

@@ 3. Main theorem: generic case
Let $A$ be a constant $2\times 2$ matrix and
$g_{z}(x)=\frac{A}{x-z},\quad z\in{\mathbb{C}}\setminus{\mathsf{E}}.$
Then
$({\mathcal{M}}_{\varkappa}g_{T^{\circ k}(0)})(x)=\frac{\Phi_{\varkappa}A\Phi_{{\varkappa}}^{*}}{-T^{\circ k}(0)(x-T(0))}+\frac{T^{\circ k}(0){\mathfrak{p}}_{\varkappa}(T^{\circ k}(0))A{\mathfrak{p}}_{\varkappa}(T^{\circ k}(0))^{*}}{x-T^{\circ(k+1)}(0)}.$ (3.1)
@@
Let us apply ${\mathcal{M}}_{\varkappa}$ to $g_{z}(x)$. To this end we consider an identity between rational functions in $z$,
$\frac{1}{2}{\mathfrak{p}}_{{\varkappa}}(z)\frac{AT^{\prime}(z)}{x-T(z)}{\mathfrak{p}}^{*}_{{\varkappa}}(z)=({\mathcal{M}}_{\varkappa}g_{z})(x)+\frac{B_{0}(x)}{z},$ (3.3)
@@
$B_{0}(x)=\frac{\Phi_{\varkappa}A\Phi^{*}_{\varkappa}}{x-T(0)},$
verified by comparing principal parts, since both sides decay at $\infty$. Set $z=T^{\circ k}(0)$ and we obtain (3.1). ∎
Define the function $w^{n}_{x}(z)$ in $z$, $n\geq 0$, by
$w^{n}_{x}(z):=\frac{1}{2^{n}}\frac{(T^{\circ n})^{\prime}(z)}{x-T^{\circ n}(z)}=\frac{zT(z)\cdots T^{\circ(n-1)}(z)}{{x-T^{\circ n}(z)}},\quad x\in[-\xi,\xi],$ (3.4)
where by definition $T^{\circ 0}(z)=z$. Note that $w_{x}^{n}$ has the partial fraction decomposition
$w_{x}^{n}(z)=\frac{1}{2^{n}}\sum_{T^{n}(y)=x}\frac{1}{y-z},$ (3.5)
in particular $w_{x}^{n}$ is a Herglotz function.
@@
We can rewrite (3.1) in terms of $w^{k}_{x}(T(0))$ as follows
@@
Let ${\mathfrak{M}}_{n,{\varkappa}}={\mathcal{M}}_{\hat{\underline{s}}^{n-1}{\varkappa}}\cdots{\mathcal{M}}_{{\varkappa}},$ so that
$({\mathfrak{M}}_{n,{\varkappa}}^{*}\Sigma_{\hat{}\underline{s}^{n}{\varkappa}})(dx)=\Sigma_{{\varkappa}}(dx).$
@@
$h_{n}(x)=({\mathfrak{M}}_{n,{\varkappa}}h_{0})(x),\quad h_{0}(x)=w_{x}^{0}(T(0))I.$
Then the sequence $h_{n}(x)$ is uniformly bounded on $\xi,\xi]$.
@@
Acting one by one by the operators ${\mathcal{M}}_{\hat{\underline{s}}^{k}{\varkappa}}$ and applying (3.2) we obtain that $h_{n}(x)$ is of the form
$h_{n}(x)=\sum_{k=0}^{n}w^{k}_{x}(T(0))A_{k}^{n},$ (3.7)
@@
From (3.5) we obtain
$w^{k}_{x}(T(0))=\frac{1}{2^{k}}\sum_{T^{\circ k}(y)=x}\frac{1}{y+\lambda}$
and since $\{y:T^{\circ k}(y)=x\}\subset[-\xi,\xi]$, we have
@@
$\int\text{\rm tr}(h_{n}(x)\Sigma_{\hat{}\underline{s}^{n}{\varkappa}}(dx))=\int\text{\rm tr}(({\mathfrak{M}}_{n,{\varkappa}}h_{0})(x)\Sigma_{\hat{}\underline{s}^{n}{\varkappa}}(dx))=\int\text{\rm tr}(h_{0}(x)\Sigma_{\varkappa}(dx))$
@@
$\int\text{\rm tr}(h_{0}(x)\Sigma_{\varkappa}(dx))=\int\text{\rm tr}(h_{n}(x)\Sigma_{\hat{}\underline{s}^{n}{\varkappa}}(dx))\geq\frac{1}{\lambda+\xi}\sum_{k=0}^{n}\text{\rm tr}(A^{n}_{k}\int\Sigma_{\hat{}\underline{s}^{n}{\varkappa}}(dx))$
@@
$\int\Sigma_{\hat{}\underline{s}^{n}{\varkappa}}(dx)=I,$
@@
$\text{\rm tr}\,h_{n}(x)\leq\frac{\lambda+\xi}{\lambda-\xi}\int\text{\rm tr}(h_{0}(x)\Sigma_{\varkappa}(dx)).$ (3.11)
Remark 3.4.
To have estimation for $\text{\rm tr}\,h_{n}(x)$ from below we write
$\text{\rm tr}\,h_{n}(x)\geq\frac{1}{\lambda+\xi}\sum_{k=0}^{n}\text{\rm tr}A_{k}^{n}.$ (3.12)
Also
$\int\text{\rm tr}(h_{0}(x)\Sigma_{\varkappa}(dx))=\int\text{\rm tr}(h_{n}(x)\Sigma_{\hat{}\underline{s}^{n}{\varkappa}}(dx))\leq\frac{1}{\lambda-\xi}\sum_{k=0}^{n}\text{\rm tr}A^{n}_{k}$
In combination with (3.12)
$\text{\rm tr}\,h_{n}(x)\geq\frac{\lambda-\xi}{\lambda+\xi}\int\text{\rm tr}(h_{0}(x)\Sigma_{\varkappa}(dx)).$ (3.13)
However it does not mean that the matrices $h_{n}(x)$ themselves are uniformly bounded from below (i.e., $h_{n}(x)\geq\delta\cdot I$, $\delta>0$). For special ${\varkappa}\in{\mathcal{F}}_{N}$ such boundedness is shown in Lemma 4.4.
Theorem 3.5.
If ${\varkappa}\in{\mathcal{G}}_{\delta}$, then the spectrum of $J_{\varkappa}$ is singular continuous.
Proof.
Assume that for some $x_{0}\in{\mathsf{E}}$,
@@
Since ${\varkappa}\in{\mathcal{G}}_{\delta}$, we also have that ${\varkappa}\in{\mathcal{G}}_{\delta}$, so the dyadic expansion of $\varkappa$} contains arbitrarily long series of consecutive zeros or ones. Using (2.10), in the first case there is a subsequence $I_{\varkappa}\subset{\mathbb{Z}}_{+}$ such that
$\lim_{n\in I_{\varkappa}}\hat{}\underline{s}^{n}{\varkappa}=0$
(we use this notation for the limit along the subsequence $I_{\varkappa}$). Alternatively, there is $I_{\varkappa}$ such that
$\lim_{n\in I_{\varkappa}}\hat{}\underline{s}^{n}{\varkappa}=1.$
@@
$\int g(x)\text{\rm tr}\left((h_{n}(x)\Sigma_{\hat{}\underline{s}^{n}{\varkappa}}(dx)\right)=$ $\int g(x)\text{\rm tr}\left(({\mathfrak{M}}_{\hat{}\underline{s}^{n}{\varkappa}}h_{0})(x)\Sigma_{\hat{}\underline{s}^{n}{\varkappa}}(dx)\right)$
$=$ $\int\text{\rm tr}\left(({\mathfrak{M}}_{\hat{}\underline{s}^{n}{\varkappa}}((g\circ T^{\circ n})h_{0}))(x)\Sigma_{\hat{}\underline{s}^{n}{\varkappa}}(dx)\right)$
$=$ $\int\text{\rm tr}\left(g(T^{\circ n}(x))h_{0}(x)({\mathfrak{M}}_{\hat{}\underline{s}^{n}{\varkappa}}^{*}\Sigma_{\hat{}\underline{s}^{n}{\varkappa}})(dx)\right)$
@@
We choose a subsequence $I_{{\varkappa},x_{0}}\subset I_{\varkappa}$ such that
$\lim_{\begin{subarray}{c}n\in I_{{\varkappa},x_{0}}\end{subarray}}T^{\circ n}(x_{0})=y_{0}.$
Now, let us fix an arbitrary small $\delta$-vicinity of $y_{0}$, $V_{\delta}=({y_{0}-\delta},{y_{0}+\delta})$. For sufficiently large $n\in I_{{\varkappa},x_{0}}$ we have $T^{\circ n}(x_{0})\in V_{\delta}$ and therefore by (3)
$\int_{V_{\delta}}\text{\rm tr}(h_{n}(x)\Sigma_{\hat{}\underline{s}^{n}{\varkappa}}(dx))=\int_{(T^{\circ n})^{-1}V_{\delta}}\text{\rm tr}(h_{0}(x)\Sigma_{{\varkappa}}(dx))\geq\text{\rm tr}(h_{0}(x_{0})\Sigma_{{\varkappa}}(\{x_{0}\}))$
@@
$C\text{\rm tr}\Sigma_{\hat{}\underline{s}^{n}{\varkappa}}(V_{\delta})\geq\text{\rm tr}(h_{0}(x_{0})\Sigma_{{\varkappa}}(\{x_{0}\}))$
with a uniform constant $C$. Passing to the limit in $n$, by continuity in ${\mathbb{Z}}_{2}$, we get for some value of ${\epsilon}\in\{0,1\}$,
$C\text{\rm tr}\Sigma_{{\epsilon}}(V_{\delta})\geq\text{\rm tr}(h_{0}(x_{0})\Sigma_{{\varkappa}}(\{x_{0}\})).$
Since $\delta$ is arbitrarily small, the spectral matrix measure $\Sigma_{{\epsilon}}$ associated to the Jacobi matrix $J_{\epsilon}$ has a non-trivial mass at $y_{0}$. Since the spectrum of both matrices $J_{0}$ and $J_{1}$ is singular continuous, we arrive to a contradiction. Thus, the initial assumption (3.14) is false, that is, $\Sigma_{{\varkappa}}(\{x_{0}\})=0$. ∎
4. Main theorem: special cases
4.1. More on the test functions $h_{n}$
Recall
${\mathfrak{M}}_{n,{\varkappa}}={\mathcal{M}}_{\hat{\underline{s}}^{n-1}{\varkappa}}\cdots{\mathcal{M}}_{{\varkappa}},\quad n\geq 1,$
so that
$({\mathfrak{M}}_{n,{\varkappa}}^{*}\Sigma_{\hat{}\underline{s}^{n}{\varkappa}})(dx)=\Sigma_{{\varkappa}}(dx).$
Let us define
${\mathfrak{P}}^{n+1}_{{\varkappa}}(z)={\mathfrak{p}}_{\hat{}\underline{s}^{n}{\varkappa}}(T^{\circ n}(z)){\mathfrak{P}}^{n}_{{\varkappa}}(z),\quad{\mathfrak{P}}^{0}_{{\varkappa}}(z)=I.$
and
$\Psi_{\varkappa}=\begin{cases}\begin{pmatrix}a_{{\varkappa}-1}\\ a_{{\varkappa}}\end{pmatrix},\quad{\varkappa}=2\underline{s}{\varkappa},\\ \\ \begin{pmatrix}a_{{\varkappa}}\\ a_{{\varkappa}+1}\end{pmatrix},\quad{\varkappa}=1+2\underline{s}{\varkappa}.\end{cases}\quad\Upsilon_{\varkappa}^{*}=\begin{cases}\begin{pmatrix}1&0\end{pmatrix},\quad{\varkappa}=2\underline{s}{\varkappa},\\ \\ \begin{pmatrix}0&1\end{pmatrix},\quad{\varkappa}=1+2\underline{s}{\varkappa},\end{cases}$
so that, see (3.2), $\Phi_{\varkappa}=\Psi_{\varkappa}\Upsilon_{\varkappa}^{*}.$ For two-dimensional vectors and matrices $A$ we define $A^{\star 2}=AA^{*}$.
Lemma 4.1.
Let $h_{n}(x)=({\mathfrak{M}}_{n,{\varkappa}}h_{0})(x)$ and $h_{0}(x)=w_{x}^{0}(T(0))I$. There are constants $\{f^{k}_{0}\}_{k=1}^{n}$ (do not depend on $x$), $f_{0}^{k}=f_{0}^{k}({\varkappa})$, such that
$h_{n}(x)=$ $w^{n}_{x}(T(0)){\mathfrak{P}}^{n}_{{\varkappa}}(T(0)){\mathfrak{P}}^{n}_{{\varkappa}}(T(0))^{*}$
$+$ $\sum_{i=1}^{n}f^{i}_{0}({\varkappa})w^{n-i}_{x}(T(0))({\mathfrak{P}}^{n-i}_{\hat{}\underline{s}^{i}{\varkappa}}(T(0))\Psi_{\hat{}\underline{s}^{i-1}{\varkappa}})^{\star 2}$ (4.1)
Moreover, let
${\mathfrak{L}}_{n,i}({\varkappa})=w^{n-1-i}_{0}(T(0))\langle{\mathfrak{P}}^{n-1-i}_{\hat{}\underline{s}^{i}{\varkappa}}(T(0))\Psi_{\hat{}\underline{s}^{i-1}{\varkappa}},\Upsilon_{\hat{}\underline{s}^{n-1}{\varkappa}}\rangle^{2},\quad i<n.$ (4.2)
Then the following recurrence relation holds
$f^{n}_{0}({\varkappa})=\sum_{i=1}^{n-1}{\mathfrak{L}}_{n,i}f^{i}_{0}({\varkappa})+w_{0}^{n-1}(T(0))\langle({\mathfrak{P}}^{n-1}_{{\varkappa}}(T(0)))^{\star 2}\Upsilon_{\hat{\underline{s}}^{n-1}{\varkappa}},\Upsilon_{\hat{\underline{s}}^{n-1}{\varkappa}}\rangle$ (4.3)
@@ 4.1. More on the test functions $h_{n}$
According to (3.2), we have
$h_{1}(x)=({\mathcal{M}}_{\varkappa}h_{0})(x)=f^{1}_{0}w^{0}_{x}(T(0))\Psi_{\varkappa}\Psi_{\varkappa}^{*}+w^{1}_{x}(T(0)){\mathfrak{p}}_{\varkappa}(T(0))^{\star 2},$
with $f^{1}_{0}=w_{0}^{0}(T(0))\Upsilon^{*}_{\varkappa}\Upsilon_{\varkappa}=w_{0}^{0}(T(0))$. Assuming (4.1) and using (3.2), we get
${\mathcal{M}}_{\hat{}\underline{s}^{n}{\varkappa}}h_{n}(x)$ $=w_{x}^{n+1}(T(0)){\mathfrak{P}}^{n+1}_{{\varkappa}}(T(0))^{\star 2}$
$+$ $w_{0}^{n}(T(0))w_{x}^{0}(T(0))(\Phi_{\hat{}\underline{s}^{n}{\varkappa}}{\mathfrak{P}}^{n}_{{\varkappa}}(T(0)))^{\star 2}$
$+$ $\sum_{i=1}^{n}f_{0}^{i}w_{0}^{n-i}(T(0))w_{x}^{0}(T(0))\left(\Phi_{\hat{}\underline{s}^{n}{\varkappa}}{\mathfrak{P}}^{n-i}_{\hat{}\underline{s}^{i}{\varkappa}}(T(0))\Psi_{\hat{}\underline{s}^{i-1}{\varkappa}}\right)^{\star 2}$
$+$ $\sum_{i=1}^{n}f_{0}^{i}w_{x}^{n-i+1}(T(0))\left({\mathfrak{p}}_{\hat{}\underline{s}^{n}{\varkappa}}(T^{\circ(n-i+1)}(0)){\mathfrak{P}}^{n-i}_{\hat{}\underline{s}^{i}{\varkappa}}(T(0))\Psi_{\hat{}\underline{s}^{i-1}{\varkappa}}\right)^{\star 2}$
By definition
${\mathfrak{p}}_{\hat{}\underline{s}^{n}{\varkappa}}(T^{\circ(n-i+1)}(0)){\mathfrak{P}}^{n-i}_{\hat{}\underline{s}^{i}{\varkappa}}(T(0))={\mathfrak{P}}^{n-i+1}_{\hat{}\underline{s}^{i}{\varkappa}}(T(0)).$
Thus, defining, see (4.2),
$f_{0}^{n+1}=w_{0}^{n}(T(0))\Upsilon^{*}_{\hat{}\underline{s}^{n}{\varkappa}}({\mathfrak{P}}^{n}_{{\varkappa}}(T(0)))^{\star 2}\Upsilon_{\hat{}\underline{s}^{n}{\varkappa}}$
$+\sum_{i=1}^{n}f_{0}^{i}w_{0}^{n-i}(T(0))\Upsilon^{*}_{\hat{}\underline{s}^{n}{\varkappa}}({\mathfrak{P}}^{n-i}_{\hat{}\underline{s}^{i}{\varkappa}}(T(0))\Psi_{\hat{}\underline{s}^{i-1}{\varkappa}})^{\star 2}\Upsilon_{\hat{}\underline{s}^{n}{\varkappa}}$
$=\sum_{i=1}^{n}{\mathfrak{L}}_{n+1,i}f^{i}_{0}({\varkappa})+w_{0}^{n}(T(0))\langle({\mathfrak{P}}^{n}_{{\varkappa}}(T(0)))^{\star 2}\Upsilon_{\hat{\underline{s}}^{n}{\varkappa}},\Upsilon_{\hat{\underline{s}}^{n}{\varkappa}}\rangle$
(4.1) and (4.3) follow by induction. ∎
4.2. Equivalence Lemma
Lemma 4.2.
Let ${\varkappa}\in{\mathcal{F}}_{N}$. The following statements are equivalent
  • (i)
    Test functions $h_{n}(x)$ are uniformly bounded from below on $\xi,\xi]$, i.e., there exists $C_{1}>0$ such that $h_{n}(x)\geq C_{1}I$, $x\in[-\xi,\xi]$.
  • (ii)
    The recurrence coefficients $f^{n}_{0}({\varkappa})$ are uniformly bounded from below, i.e., there exists $C_{2}>0$ such that $f^{n}_{0}({\varkappa})\geq C_{2}$.